OpenStudy (anonymous):

Tan(INVsin3/4- INVcos1/5)

OpenStudy (jdoe0001):

\(\bf tan\left[\cfrac{sin^{-1}\left(\frac{3}{4}\right)}{cos^{-1}\left(\frac{1}{5}\right)}\right] \quad ? \)

OpenStudy (anonymous):

It's tan(sin-cos)

OpenStudy (jdoe0001):

ohh yea.. so \(\bf tan\left[sin^{-1}\left(\frac{3}{4}\right)-cos^{-1}\left(\frac{1}{5}\right)\right]\)

OpenStudy (jdoe0001):

keep in mind that the inverse trig functions, for sine, will have a range between \(\bf \cfrac{\pi}{2}\textit{and }-\cfrac{\pi}{2}\) and the inverse function for cosine has a range of \(\bf 0 \ to \pi\)

OpenStudy (anonymous):

 after drawing the triangles and plugging it to the formula is this correct?